# Deep learning-based sequential data assimilation for chaotic dynamics identifies local instabilities from single state forecasts

### v1

@@ IV.3 Numerical results
The surrogate $\mathbf{P}^{\mathrm{a}}$, denoted $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and estimated from Eq. (11), is compared to that of a concurrently running well-tuned EnKF with $N_{\mathrm{e}}=40$, whose analysis error covariance matrix is $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ is compared to $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ using a normalized Bures-Wasserstein distance: (Bhatia, Jain, and Lim, 2019)
@@
where $\mathbf{U}$ and $\mathbf{V}$ are two semi-definite symmetric matrices. This metric is expected to smoothly account for the unmatched principal axes of $\mathbf{U}$ and $\mathbf{V}$, but also their associated variances (eigenspectra). The time-averaged $d_{\mathrm{BW}}$ distance between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ is $0.013$ whereas it is $0.048$ between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $(0.40)^{2}\mathbf{I}_{\mathrm{x}}$, which approximates $\mathbf{P}^{\mathrm{a}}$ of a well-tuned 3D-Var. Furthermore, the averaged-in-time eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are plotted in Fig. 1. They are remarkably close to each other for the first $10$ modes. Beyond these modes the $a_{\bm{\theta}}$ operator is likely to selectively apply some multiplicative inflation, as one would expect from such stable DA runs.
@@
We further compute the principal angles (Bocquet and Carrassi, 2017) of the vector subspaces generated by the $N_{\mathrm{u}}=14$ dominant eigenvalues of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and of $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. Recall that $N_{\mathrm{u}}=14$ is the dimension of the L96 ${\mathcal{U}}$. The principal angles are intrinsic to the relative position of these subspaces; they do not depend on any coordinate system used to parameterize them. This indicates how close the most unstable directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are in state space. The principal angles are reported in Fig. 2. Overall, the simplex formed by the EnKF is on average the most aligned with ${\mathcal{U}}$.(Bocquet and Carrassi, 2017) The subspace spanned by the dominant axes of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ are also well aligned with ${\mathcal{U}}$, yet progressively diverges when incorporating less unstable directions. For comparison, the principal angles of ${\mathcal{U}}$ with an isotropically randomly sampled $N_{\mathrm{u}}=14$-dimensional subspace are also computed.
@@ IV.4 Main interpretation
These numerical results indicate that $a_{\bm{\theta}}$ defined through Eq. (2a), depends on the innovation, but also on the single forecast state when $N_{\mathrm{e}}=1$. This does not hold for the EnKF incremental update which only indirectly depends on the forecast state via the ensemble-based forecast error covariances. Hence, without the need of an ensemble, $a_{\bm{\theta}}$ extracts from the forecast state critical pieces of information on the unstable directions, as shown by the principal angles experiment.
@@ IV.5 Locality and scalability
Next, we have trained $a_{\bm{\theta}}$ on the L96 model, using the reference configuration, but with a changing state space dimension $N_{\mathrm{x}}$ in between $20$ and $160$. The aRMSEs of well-tuned EnKFs for the changing $N_{\mathrm{x}}$, and picking $N_{\mathrm{e}}=N_{\mathrm{x}}$, has been computed for comparison. The test DAN aRMSEs shows no significant dependence on $N_{\mathrm{x}}$ and are all within $5\%$ of the EnKFs. Hence, because the performance of $a_{\bm{\theta}}$ with an unchanged architecture and number of parameters is barely affected by increasing $N_{\mathrm{x}}$, we conjecture that the learned analysis extracts local pieces of information from the forecast state.
If true, the $a_{\bm{\theta}}$ operator learned for DA on an $N_{\mathrm{x}}=40$ L96 model could be applied directly to an L96 DA run with a different $N_{\mathrm{x}}$. Recall that the L96 states typically exhibit local highs and lows of Rossby-like waves, whose number scales linearly with $N_{\mathrm{x}}$. Thus as long as the spatial extent of those waves is captured by the receptive field of the CNN, the same layers of $a_{\bm{\theta}}$ with the same weights and biases might be able to handle L96 states of distinct dimensionality.
To test this hypothesis, we use the same $a_{\bm{\theta}}$ operator (same weights and biases) learned as before with $N_{\mathrm{e}}=1$, $N_{\mathrm{x}}=40$ but apply it to L96 models with $N_{\mathrm{x}}$ ranging from $20$ to $160$. The corresponding aRMSEs are reported in Fig. 3. This transdimensional transfer works surprisingly well: the aRMSEs are roughly the same for all $N_{\mathrm{x}}$ (between 0.188 and 0.197).
Figure 3: Test aRMSEs of $a_{\bm{\theta}}$ operators learned from either L96 models with varying $N_{\mathrm{x}}$, or learned from the $N_{\mathrm{x}}=40$ L96 model but applied to varying $N_{\mathrm{x}}$ L96 models. The dimension $N_{\mathrm{u}}$ of ${\mathcal{U}}$ is much steeper compared to the slowly increasing aRMSE curves.
This strongly supports the fact that $a_{\bm{\theta}}$ extracts local information from the forecast state (of various dimension in this experiment), relying on its convolution layers. It is therefore able to capture where, in phase-space, the error mass is concentrated. These localized error structures should be related to the localization of the dominant covariant Lyapunov vectors.(Pazó, Rodríguez, and López, 2010; Vannitsem and Lucarini, 2016; Giggins and Gottwald, 2019)
@@ V Conclusions
We have demonstrated on the L96 chaotic model that a learned DL-based analysis $a_{\bm{\theta}}$, key part of a sequential DA scheme, can be almost as accurate as the best possibly tuned EnKF, the benchmark for ensemble filtering methods in this model. More importantly, this learned DA scheme does not require any ensemble and can equally well rely on a single state forecast. Therefore, $a_{\bm{\theta}}$ can presumably read local patterns, that are representative of unstable and uncertain modes, from the forecast state alone. We believe that this is fundamentally made possible by some multiplicative ergodic theorem applied to sequential DA seen as a non-autonomous random dynamical system driven by time-dependent true dynamics and observation operators, and white-in-time observation errors.
@@
What is achieved by $a_{\bm{\theta}}$ resonates with the parametric EnKF, (Pannekoucke et al., 2016; Pannekoucke, Bocquet, and Ménard, 2018) which encodes the errors of the day in a couple of dynamical ancillary fields, preventing the use of an ensemble. Amazingly, the learned $a_{\bm{\theta}}$ exhibited here is even more radical and extracts that information from the state itself.
Taking a step back, we learned from DL that an accurate and efficient DA analysis operator could capture the dynamical error without an ensemble, leveraging model-specific information. This argues in favor of rethinking popular sequential DA schemes for chaotic dynamics.
@@ Acknowledgements.
The authors acknowledge the support of the project SASIP (grant no. $353$) funded by Schmidt Sciences – a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies. CEREA is a member of Institut Pierre-Simon Laplace (IPSL).
@@
Appendix A Architecture of the analysis operator
The learned operator $a_{\bm{\theta}}$ predicts the analysis increment to be added to the forecast/background ensemble, following Eq. (2) The architecture of the network is a basic residual CNN. It begins with an initial convolution that takes $N_{\mathrm{e}}+1$ channels as inputs and, with $N_{\mathrm{f}}$ filters, outputs $N_{\mathrm{f}}$ channels. This initial layer is followed by $N_{\mathrm{b}}$ residual blocks. Each one of these blocks is a succession of $N_{\mathrm{sb}}$ sub-blocks. Each subblock is made of: (i) a convolutional layer with $N_{\mathrm{f}}$ channels as inputs, has $N_{\mathrm{f}}$ filters and a kernel size for each of its filter of $f_{\mathrm{v}}$, (ii) a batch normalization layer, and (iii) an activation function chosen to be mish. (Misra, 2019) The CNN finishes with a final convolutional layer that takes $N_{\mathrm{f}}$ channels as inputs and, with $N_{\mathrm{e}}$ filters, outputs $N_{\mathrm{e}}$ channels. The kernel size of the initial and final channels is $f_{\mathrm{v}}$. A schematic of the CNN architecture is displayed in Fig. 4. The reference configuration, $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$, has about $2\times 10^{5}$ trainable parameters.
Figure 4: Architecture of the residual convolutional network, where $N_{\mathrm{b}}=2$, $N_{\mathrm{sb}}=3$. $\mathrm{conv}_{N_{1},N_{2},f}$ is a generic one-dimensional convolutional layer of dimension $N_{1}$, with $N_{2}$ filters of kernel size $f$.

### v2

@@ IV.3 What is learned? - Supporting numerical results
The surrogate $\mathbf{P}^{\mathrm{a}}$, denoted $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and estimated from Eq. (11), is compared to that of a concurrent well-tuned EnKF with $N_{\mathrm{e}}=40$, whose analysis error covariance matrix is $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ is compared to $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ using a normalized Bures-Wasserstein distance: (Bhatia, Jain, and Lim, 2019)
@@
where $\mathbf{U}$ and $\mathbf{V}$ are two semi-definite symmetric matrices. This metric is expected to smoothly account for the unmatched principal axes of $\mathbf{U}$ and $\mathbf{V}$, but also their associated variances (eigenspectra). The time-averaged $d_{\mathrm{BW}}$ distance between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ is $0.013$ whereas it is $0.048$ between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $(0.40)^{2}\mathbf{I}_{\mathrm{x}}$, which approximates $\mathbf{P}^{\mathrm{a}}$ of a well-tuned 3D-Var. The time-averaged eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are plotted in Fig. 2. They are remarkably close to each other for the first $10$ modes. Beyond these modes the $a_{\bm{\theta}}$ operator is likely to selectively apply some (multiplicative) inflation, as one would expect from such stable DA runs.
@@
We further compute the principal angles(Bocquet and Carrassi, 2017) of the vector subspaces generated by the $N_{\mathrm{u}}=14$ dominant eigenvalues of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and of $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. They are reported in Fig. 3. Recall that $N_{\mathrm{u}}=14$ is the dimension of the L96 ${\mathcal{U}}$. The principal angles are intrinsic to the relative position of these subspaces; they do not depend on any coordinate system used to parameterize them. This indicates how close the most unstable directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are in state space. From Fig. 3, we observe that the simplex formed by the EnKF is on average the most aligned with ${\mathcal{U}}$.(Bocquet and Carrassi, 2017) The subspace spanned by the dominant axes of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ are also well aligned with ${\mathcal{U}}$, yet progressively diverges when incorporating less unstable directions. For comparison, the principal angles of ${\mathcal{U}}$ with an isotropically randomly sampled $N_{\mathrm{u}}=14$-dimensional subspace are also plotted in Fig. 3.
@@ IV.4 Main interpretation
These numerical results indicate that $a_{\bm{\theta}}$ defined through Eq. (2a), depends on the innovation, but also on the single forecast state when $N_{\mathrm{e}}=1$. This does not hold for the EnKF incremental update which only indirectly depends on the forecast state via the ensemble-based forecast error covariances. Hence, without the need for an ensemble, $a_{\bm{\theta}}$ extracts from the forecast state critical pieces of information on the unstable directions, as shown by the principal angles experiment.
@@ IV.5 Locality and scalability
Next, we have trained $a_{\bm{\theta}}$ on the L96 model using the reference configuration with $N_{\mathrm{e}}=1$, but with a changing state space dimension $N_{\mathrm{x}}$ in between $20$ and $160$. The aRMSEs of well-tuned EnKFs for the changing $N_{\mathrm{x}}$, and picking $N_{\mathrm{e}}=N_{\mathrm{x}}$, has been computed for comparison. The test DAN aRMSEs shows no significant dependence on $N_{\mathrm{x}}$ and are all within $5\%$ of the EnKFs. Hence, because the performance of $a_{\bm{\theta}}$ with an unchanged architecture and the same number of parameters is barely affected by increasing $N_{\mathrm{x}}$, we conjecture that the learned analysis extracts local pieces of information from the forecast state.
If true, the $a_{\bm{\theta}}$ operator learned for DA on an $N_{\mathrm{x}}=40$ L96 model could be applied directly to an L96 DA run with a different $N_{\mathrm{x}}$. Recall that the L96 states exhibit local highs and lows of Rossby-like waves, whose number scales linearly with $N_{\mathrm{x}}$. Thus as long as the spatial extent of those waves is captured by the receptive field of the CNN, the same layers of $a_{\bm{\theta}}$ with the same weights and biases might be able to handle L96 states of distinct dimensionality.
To test this hypothesis, we use the same $a_{\bm{\theta}}$ operator (same weights and biases) learned as before with $N_{\mathrm{e}}=1$, $N_{\mathrm{x}}=40$ but apply it now to L96 models with $N_{\mathrm{x}}$ ranging from $20$ to $160$. The corresponding aRMSEs are reported in Fig. 4, which shows that these aRMSEs are roughly the same for all $N_{\mathrm{x}}$ (between 0.188 and 0.197). This demonstrates that this transdimensional transfer works surprisingly well.
Figure 4: Test aRMSEs (blue and yellow full lines) of $a_{\bm{\theta}}$ operators learned from either L96 models with varying $N_{\mathrm{x}}$, or learned from the $N_{\mathrm{x}}=40$ L96 model but applied to varying $N_{\mathrm{x}}$ L96 models. The dimension $N_{\mathrm{u}}$ of ${\mathcal{U}}$ (gray dashed line) is much steeper compared to the slowly increasing aRMSE curves.
This strongly supports the fact that $a_{\bm{\theta}}$ extracts local information from the forecast state (of various dimensions in this experiment), relying on its convolution layers. It is therefore able to capture where, in phase-space, the error mass is concentrated. We hypothesize that these localized error structures are related to the localization of the dominant covariant Lyapunov vectors.(Pazó, Rodríguez, and López, 2010; Vannitsem and Lucarini, 2016; Giggins and Gottwald, 2019) A proper mathematical definition of such spatial localization can be found in these references.
@@ V Conclusions
Using the L96 chaotic model we have demonstrated that a learned DL-based analysis $a_{\bm{\theta}}$, key part of a sequential DA (often referred to as a filtering scheme) can be almost as accurate as the best possibly tuned EnKF, the benchmark for ensemble filtering methods in this model. More importantly, this learned DA scheme does not require any ensemble and can equally well rely on a single state forecast. Therefore, $a_{\bm{\theta}}$ appears to be able to retrieve local patterns, representative of unstable and uncertain modes, from the forecast state alone. We believe that this is fundamentally made possible by some multiplicative ergodic theorem applied to sequential DA seen as a non-autonomous random dynamical system driven by time-dependent true dynamics and observation operators, and white-in-time observation errors.
@@
What is achieved by $a_{\bm{\theta}}$ resonates with the parametric EnKF, (Pannekoucke et al., 2016; Pannekoucke, Bocquet, and Ménard, 2018) which encodes the errors of the day in a couple of dynamical ancillary fields, preventing the use of an ensemble. Amazingly, our learned $a_{\bm{\theta}}$ is even more radical and extracts that information from the state itself.
Taking a step back, we learned from DL that an accurate and efficient DA analysis operator could capture the dynamical error without an ensemble, leveraging model-specific information. This promotes a rethinking of the popular sequential DA schemes for chaotic dynamics.
Supplementary material
See the supplementary material for further details on the evaluation of the $a_{\bm{\theta}}$-based DA method, as reported in Sec. III. Specifically, a table of the key aRMSE scores mentioned Sec. III is provided, as well as plots of the aRMSE curves related to the sensitivity experiments mentioned in Sec. III.
@@ Acknowledgements.
The authors are grateful to the Editor Jürgen Kurths and an anonymous Reviewer for their comments and suggestions on the original version of the manuscript. The authors acknowledge the support of the project SASIP (grant no. $G-24-66154$) funded by Schmidt Sciences – a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies. CEREA is a member of Institut Pierre-Simon Laplace (IPSL).
@@
Supplementary Material: Accurate deep learning-based filtering for chaotic dynamics by identifying instabilities without an ensemble
In the following DA, DL, CNN, EnKF, RMSE and aRMSE are the initialisms and acronyms of data assimilation, deep learning, convolutional neural network, ensemble Kalman filter, root mean square error (of the analysis state versus the true state), and time-averaged root mean square error (of the analysis state versus the true state), respectively. All the $a_{\bm{\theta}}$ operators trained and tested here are based on $N_{\mathrm{e}}=1$, i.e. no ensemble but a single state forecast.
V.1 Impact of the training dataset size on the DA tests
A series of $a_{\bm{\theta}}$ operators are learned from increasingly larger training and validation datasets (per epoch), which is achieved by increasing the number of trajectories $N_{\mathrm{r}}$ in the datasets in the range $[2^{10},2^{22}]$, and displayed in Fig. 5. The hyperparameters of the CNN architecture are those of the reference configuration.
Figure 5: Test aRMSEs of $a_{\bm{\theta}}$ operators learned from a L96 models with a varying number of trajectories $N_{\mathrm{r}}$ in the training dataset, commensurate to the size of the training dataset per epoch.
The batch size, i.e. the number of trajectories in the training batch, is chosen in this experiment to be $\mathrm{min}\left(N_{\mathrm{r}}/10,2^{11}=2,048\right)$ in order to be able to maintain the targeted training and validation datasets ratio, while seeking a large batch for numerical efficiency. As a compromise between accuracy and computational time for this study, most of the experiments are conducted with $N_{\mathrm{r}}=2^{18}=262,144$.
V.2 Impact of the number of filters of $a_{\bm{\theta}}$ as a CNN on the DA tests
A series of $a_{\bm{\theta}}$ operators are learned from an increasingly larger number of channels/filters $N_{\mathrm{f}}$. All other hyperparameters of the CNN architecture are those of the reference configuration.
Figure 6: Test aRMSEs of the DAN method applied to the L96 model, as a function of the number of channels/filters $N_{\mathrm{f}}$ in $a_{\bm{\theta}}$.
As a compromise between accuracy and computational time for this paper, most of the experiments are conducted with $N_{\mathrm{f}}=40$.
V.3 Impact of the observation noise magnitude on the DA tests
The impact of observation noise in the DA test experiments with magnitude distinct from the noise magnitude used in the training of $a_{\bm{\theta}}$ is discussed in the paper in Section III.B. Figure 7 displays the corresponding DAN and EnKF aRMSE curves. In all cases the analysis aRMSE remains below the corresponding observational error standard deviation $\sigma_{\mathrm{y}}$.
Figure 7: Test aRMSEs of the DAN method, and of the EnKF $N_{\mathrm{e}}=20,40$ applied to the L96 model, as a function of the true standard deviation of the assimilated observations.
V.4 Impact of the sparsity of the observations on the DA tests
The impact on the test aRMSEs of the DAN scheme with non-trivial observation operators $\mathcal{H}_{k}$ is discussed in the paper in Section III.B, with an experiment where $a_{\bm{\theta}}$ is trained with varying $\mathcal{H}_{k}$ of time-dependent random sparsity ratios, and then tested with observation operators of fixed in time sparsity ratio. Figure 8 displays the corresponding DAN and EnKF aRMSE curves. The observation error standard deviation is $\sigma_{\mathrm{y}}=1$.
Figure 8: Test aRMSEs of the DAN method, and of the EnKF $N_{\mathrm{e}}=20,40$ applied to the L96 model, as a function of the observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$.
Moreover, we report here (though not in the paper) that, as a sanity check, we learned $a_{\bm{\theta}}$ and tested the resulting DAN scheme with a staggered observation operator, where only the variables of the odd sites of the L96 model are observed (a classical test in methodological DA studies). We obtain for the DL-based DA scheme and a well-tuned EnKF with $N_{\mathrm{e}}=40$, aRMSEs of $0.289$ and $0.288$, respectively, which suggests that non-trivial observation operators are properly dealt with by $a_{\bm{\theta}}$.
V.5 Key aRMSE scores of the main DA methods
For comparison and for reference, the aRMSE scores of the classical DA methods and of the DL-based DAN schemes tested in this paper, are compiled in Tab. 1. The tested well-tuned EnKFs are based on the finite-size EnKF (EnKF-N) DA method with has implicit adaptive inflation and relies here on an ensemble size of either $N_{\mathrm{e}}=20$ or $N_{\mathrm{e}}=40$, beyond the unstable-neutral subspace dimension $N_{\mathrm{u}}=14$ where localization is unnecessary. The architecture and hyperparameters of the DL-based DA schemes are those of the reference configuration described in the paper, with possibly the exception of $N_{\mathrm{e}}$ and $N_{\mathrm{f}}$ which are hence explicitly specified in the table.
Table 1: Compilation of the key aRMSE scores of the classical and DL-based DA methods tested in this paper.
DA method well-tuned classical DL-based aRMSE
EnKF-N, $N_{\mathrm{e}}=20$ yes $0.191$
EnKF-N, $N_{\mathrm{e}}=40$ yes $0.179$
3D-Var yes $0.40$
$a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.191$
$a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=100$ yes $0.185$
linear $a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.384$
simplified ${\hat{a}}_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.382$
V.6 Typical RMSE time series of DAN and of a well-tuned EnKF
Figure 9 illustrates time series of instant RMSEs for the DAN scheme in the reference configuration with $N_{\mathrm{e}}=1$, and for a well-tuned EnKF with $N_{\mathrm{e}}=40$, actually the EnKF-N. These time series include the initial spin-up time period. Note that the EnKF starts from a perturbed true state, whereas DAN starts from a random state on the L96 attractor.
Figure 9: Instant RMSEs of the DAN scheme ($N_{\mathrm{e}}=1$) and of a well-tuned EnKF ($N_{\mathrm{e}}=40$) as a function of the L96 model time.